# Neural Algorithmic Reasoning with Multiple Correct Solutions

### v1

@@ 1 Introduction
Classical algorithms such as Merge Sort guarantee correctness and generalize perfectly on abstract inputs [1]. Unfortunately, real problems are rarely abstract inputs. Consider pathfinding: to give traffic directions via a shortest paths algorithm, one must compress physical distance and multi-dimensional congestion data for each street into edge labels, thereby losing information. Bad edge labels mean bad directions, even despite a good algorithm. Neural Networks (NNs) avoid this problem by directly operating on high-dimensional data. Instead of compressing distance and congestion to a single path-length edge label, an NN can operate directly on distance and congestion data. The chief hurdle to NN performance is failure to generalize: common architectures are fragile to changes in input size and structure — exactly where algorithms succeed [1]. It is thus natural to wonder whether an NN might learn to generalize from an algorithm. To that end, Neural Algorithmic Reasoning trains NNs to mimic the running of classical algorithms [1]. A benefit of this approach is staying high-dimensional: no information is lost when projecting high-dimensional real data into a single abstract output for a classical algorithm [1]. Thus far, NNs have been trained to mimic algorithms which recover one solution from the space of possible correct solutions [2, 3, 4, 5, 6, 7, 8]. Ordinarily, an NN trained to perform DFS produces an array, which one can view as a distribution containing the probabilities of different solutions, from which the likeliest solution is returned as the model output. We adapt the model input to reflect multiple correct solutions and present methods of retrieving multiple solutions from the model output distribution. Using only one solution in NAR may be vulnerable to local minima in the model training. By diversifying the space of solutions returned by NAR, we may decrease the risk of solutions originating from a local minimum. For applications like cyberphysical systems, the flexibility of multiple solutions is essential. A single solution, for reasons of cost, safety, or other reasons outside the algorithmic problem, may not be amenable to a final task.
@@ 2 Methodology
To produce multiple solutions for a single graph, we train an NN to predict a distribution of solutions for each individual graph. Ordinarily, the CLRS benchmark [2] generates training data by running classical algorithms with deterministic tie-breaking. For example, the input graph in Figure 1 has two valid orders of depth-first traversal, but classical DFS will only return one based on which edge it explores first. To create our training examples, we run a classical algorithm (e.g. DFS) multiple times with randomized tiebreaking, and average the outputs (e.g. individual DFS trees) into a distribution of solutions (refer to Appendices A.1.1 and A.2.1). As we want the NN to predict a distribution, we train to minimize a distribution divergence (Kullback-Leibler: KL) [9]. Our final goal is to recover multiple solutions for an input graph (solutions being individual DFS trees or BF paths); so, we develop stochastic, algorithm-specific sampling methods. Stochasticity of sampling methods means samples are likely to have multiple different solutions. For DFS, we sample the immediate parents of vertices until we reach the root vertex (which can be seen as executing DFS in reverse on a graph); our two DFS extractors, which we call Upwards and AltUpwards, differ in the manner in which they discard potential solutions (Appendix A.1.2). For BF, Beam samples a number of parents for non-source vertices according to parent-probabilities and chooses the lowest cost path arising (Appendix A.2.2). Also for BF, Greedy samples parents for each non-source vertex and chooses the parent with the lowest-cost outgoing edge for each non-source vertex (Appendix A.2.2).
Figure 1: Blueprint for Neural Algorithmic Reasoning with multiple correct solutions. For Step 2 see Figure 3. For Step 3 see figure 4.
@@ 3 Experiments
Addressing the lack of evaluation methods for NAR with multiple solutions, we propose metrics to evaluate the validity and diversity of solutions. For a single graph, our NN predicts one solution distribution. From this distribution, we sample 5 solutions and count the number of distinct and valid solutions. For Bellman-Ford, a solution is valid if and only if it gives all shortest paths from the source vertex. For DFS, we say a solution is valid if and only if it does not contradict some necessary conditions (sufficient conditions not being known in graph theory). Refer to Figures 3 and 4 in Appendix A for illustrations of the data generation and solution extraction on a graph, respectively. We investigate the coverage of the solution space by our method in Appendix C.
We run our experiments using the CLRS Benchmark [2], which handles data generation, model training, and model evaluation in a unified way. CLRS generates graphs of size 4, 7, 11, 13, and 16 to train our models on. In keeping with established methodology [2, 3], we test our models on graphs within and outside the range of our training sizes. For within-sample testing, we choose $n=|V(G)|=5$, as solutions on these graphs are easily manually verifiable and fast to compute at scale. For our out-of-distribution evaluation, we use the default CLRS-30 benchmark testing size $n=64$. The models are trained with default CLRS-30 benchmark model parameters [2].
@@ 3.1 Results for Bellman-Ford
On small graphs, ($n=5$), sampling methods extract a single correct minimum-cost path with perfect accuracy from both the NN-predicted distributions (Table 1, Table 2). We also extract solutions from the empirical distributions in order to validate our sampling methods, as we would expect suitable sampling methods to return sensible solutions from distributions that are close to optimal. Similar results for empirical and predicted distributions suggest the NN predicts a good distribution for $n=5$ graphs. The lack of variety in solutions, with sampling methods always extracting the same solution, is expected for small graphs as there is likely only one minimum-cost path. Diversity increases in the case of $n=64$, where for each method, sampling a predecessor array $\pi$ five times gives a new solution in nearly all cases, as shown in Table 1.
For larger graphs, Table 2 shows that our sampling methods successfully yield correct minimum-cost paths through the graphs on the empirical distributions (denoted as $\mathcal{P}$), thus validating our sampling methods. However, the distributions of solutions predicted by our models, which we denote $\tilde{\mathcal{P}}$, are not sufficiently accurate to consistently have correct solutions sampled from them. These results permit us to conclude that our models perform very well on small graphs, but have difficulty scaling outside of their training distribution. While comparing our approach with the deterministic algorithmic frameworks of Veličković et al. [2] and Ibarz et al. [3] is disadvantageous to our method due to its broader scope, it is worth noting that our results are markedly worse than those in [2, 3]. However, our results are encouraging for future work, as we have demonstrated the feasibility of training neural networks to emulate a version of Bellman-Ford that is more powerful than its deterministic variant.
Table 1: The average number of unique and valid solutions obtained when sampling 5 solutions from empirical ($\mathcal{P}$) and model output ($\tilde{\mathcal{P}}$) distributions. Rounded figures reported.
$n=5$ (reported figures out of 5)
@@
Greedy (BF) $1.00\pm 0.00$ $5.00\pm 0.00$ $1.00\pm 0.00$ $5.00\pm 0.00$
Beam (BF) $1.00\pm 0.00$ $5.00\pm 0.00$ $1.00\pm 0.00$ $5.00\pm 0.00$
Upwards (DFS) $3.72\pm 0.19$ $1.58\pm 0.29$ $2.60\pm 0.23$ $0.78\pm 0.36$
AltUpwards (DFS) $2.54\pm 0.15$ $4.28\pm 0.11$ $3.66\pm 0.26$ $2.78\pm 0.69$
$n=64$ (reported figures out of 5)
@@
Greedy (BF) $5.00\pm 0.00$ $5.00\pm 0.00$ $4.98\pm 0.04$ $1.98\pm 0.94$
Beam (BF) $5.00\pm 0.00$ $5.00\pm 0.00$ $5.00\pm 0.00$ $2.72\pm 1.12$
Upwards (DFS) $5.00\pm 0.00$ $0.04\pm 0.09$ $5.00\pm 0.00$ $0.02\pm 0.04$
AltUpwards (DFS) $5.00\pm 0.00$ $0.24\pm 0.15$ $5.00\pm 0.00$ $0.06\pm 0.09$
@@ 3.2 Results for Depth-First Search
In most cases, our sampling methods perform worse than simply using the most likely predecessor for all $v\in V(G)$, but have the advantage of recovering multiple solutions (Table 1). Table 2 shows that AltUpwards sampling outperforms Upwards sampling, even being within the margin of error of Argmax for $\mathcal{P}$ in some experiments. On smaller graphs, AltUpwards performs well on the model distributions, returning a valid DFS-tree around half of the time. DFS is known to be difficult for neural algorithmic reasoning, with neural algorithmic reasoners giving correct (single) solutions only around a third of the time [2, 3]. This contextualises our results and demonstrates the promise of our approach compared to both classical algorithms and previous work in neural algorithmic reasoning.
In terms of the ability of recovering multiple solutions, Table 1 shows that for both considered graph sizes, many unique solutions are obtained, with better accuracy scores for $\mathcal{P}$ than for $\tilde{\mathcal{P}}$, particularly for large graphs, where accuracy is low in general. In fact, for the out-of-distribution testing, our NAR method for DFS performs significantly worse than the current state of the art [3]. Because our method is non-deterministic, this should not be taken as a conclusively negative result: we would expect our method for returning multiple solutions to perform worse than the results for models giving single solutions in [3], as predicting multiple solutions is a more difficult task with more noise. Furthermore, AltUpwards is generally able to recover more unique and more valid solutions than Upwards, but at a high computational cost. Table 1 shows that for large graphs, both sampling methods yield many unique solutions, but these solutions are hardly ever valid for large graphs.
Table 2: Rounded mean accuracies and standard deviations for all sampling methods over 5 runs.
@@ Conclusion
We outline a two-part approach to Neural Algorithmic Reasoning with multiple solutions: we first train models to predict distributions of solutions and second stochastically extract solutions from those distributions. We demonstrate that accurate stochastic extraction of multiple solutions from distributions of solutions is feasible, given a good distribution, but also that NNs do not easily learn to predict good distributions, which means our methods will improve with advances in NN training in this field. We attribute this difficulty to a number of factors. First, Neural Algorithmic Reasoning is difficult: previous work has shown that NNs only learn up to 30% accuracy on deterministic DFS without hint pointer reversal [4]. Second, training a model to fit a distribution of solutions does not directly optimize for correct solutions. Solutions must still be extracted from the distribution via stochastic methods, to which the loss metric does not extend (Figure 7). Improving methods for extracting solutions remains an important algorithmic task. Still, with the promise of sampling multiple solutions from a distribution demonstrated, we believe our modifications of the CLRS benchmark act as proof-of-concept for further work. While typically straightforward to apply our method to an algorithm, a downside is it requires tailored work per algorithm (especially in designing sampling methods). A more general method not requiring algorithm-specific work would be ideal, if difficult to achieve. Having multiple solutions adds essential flexibility to applications of NAR such as cyberphysical systems and might aid in explaining NAR models.
@@ Appendix A Methodology
Figure 2: Parent Tree encoding of a single DFS solution. The start node, 0, is its own parent, represented by the value 0 at index 0. The value 0 at index 1 indicates that vertex 0 is the parent of vertex 1. Similarly, the value 1 at index 2 indicates that vertex 1 is the parent of vertex 2.

### v3

@@ 1 Introduction
Classical algorithms such as Merge Sort guarantee correctness and generalize perfectly on abstract inputs [1]. Unfortunately, real problems are rarely abstract inputs. Consider pathfinding: to give traffic directions via a shortest paths algorithm, one must compress physical distance and multi-dimensional congestion data for each street into edge labels, thereby losing information. Bad edge labels mean bad directions, even despite a good algorithm. Neural Networks (NNs) avoid this problem by directly operating on high-dimensional data. Instead of compressing distance and congestion to a single path-length edge label, an NN can operate directly on distance and congestion data. The chief hurdle to NN performance is failure to generalize: common architectures are fragile to changes in input size and structure — exactly where algorithms succeed [1]. It is thus natural to wonder whether an NN might learn to generalize from an algorithm. To that end, Neural Algorithmic Reasoning trains NNs to mimic the running of classical algorithms [1]. A benefit of this approach is staying high-dimensional: no information is lost when projecting high-dimensional real data into a single abstract output for a classical algorithm [1]. Thus far, NNs have been trained to mimic algorithms which recover one solution from the space of possible correct solutions [2, 3, 4, 5, 6, 7, 8]. Ordinarily, an NN trained to perform DFS produces an array, which one can view as a distribution containing the probabilities of different solutions, from which the likeliest solution is returned as the model output. We adapt the model input to reflect multiple correct solutions and present methods of retrieving multiple solutions from the model output distribution. Using only one solution in NAR may be vulnerable to local minima in the model training. By diversifying the space of solutions returned by NAR, we may decrease the risk of solutions originating from a local minimum. For applications like cyberphysical systems, the flexibility of multiple solutions is essential. A single solution, for reasons of cost, safety, or compliance, may not be amenable to a final task.
@@ 2 Methodology
To produce multiple solutions for a single graph, we train an NN to predict a distribution of solutions for each individual graph (solutions being individual DFS trees or BF paths, their distribution represented as distributions over child-parent pairs, as discussed in A.1.1 and A.2.1). Ordinarily, the CLRS benchmark [2] generates training data by running classical algorithms with deterministic tie-breaking. For example, the input graph in Figure 1 has two valid orders of depth-first traversal, but classical DFS will only return one based on which edge it explores first. To create our training examples, we run a classical algorithm (e.g. DFS) multiple times with randomized tiebreaking, and average the outputs (e.g. individual DFS trees) into a distribution of solutions. As we want the NN to predict a distribution, we train to minimize a distribution divergence (Kullback-Leibler: KL) [9] between the input distribution of solution and that given by the NN. Our final goal is to recover multiple solutions for an input graph; so, we develop stochastic, algorithm-specific sampling methods. Stochasticity of sampling methods means samples are likely to have multiple different solutions. For DFS, we sample the immediate parents of vertices until we reach the root vertex (which can be seen as executing DFS in reverse on a graph); our two DFS extractors, which we call Upwards and AltUpwards, differ in the manner in which they discard potential solutions (Appendix A.1.2). For BF, Beam samples a number of parents for non-source vertices according to parent-probabilities and chooses the lowest cost path arising (Appendix A.2.2). Also for BF, Greedy samples parents for each non-source vertex and chooses the parent with the lowest-cost outgoing edge for each non-source vertex (Appendix A.2.2).
Figure 1: Blueprint for NAR with multiple correct solutions. For Step 2 see Figure 3. For Step 4 see Figure 4 and Appendices A.1.2 A.2.2.
@@ 3 Experiments
Addressing the lack of evaluation methods for NAR with multiple solutions, we propose metrics to evaluate the validity and diversity of solutions. For a single graph, our NN predicts one solution distribution. From this distribution, we sample 5 solutions and count the number of distinct and valid solutions. For Bellman-Ford, a solution is valid if and only if it gives all shortest paths from the source vertex. For DFS, we say a solution is valid if and only if it does not contradict some necessary conditions A.2.3 (sufficient conditions not being known from tree structure alone). Figures 3 and 4 in Appendix A illustrate data generation and solution extraction on a graph. We investigate our coverage of the solution space in Appendix C. We run our experiments using the CLRS Benchmark [2], which handles data generation, model training, and model evaluation in a unified way. CLRS generates graphs of size 4, 7, 11, 13, and 16 to train our models on. In keeping with established methodology [2, 3], we test our models on graphs within and outside the range of our training sizes. For within-sample testing, we choose $n=|V(G)|=5$, as solutions on these graphs are easily manually verifiable and fast to compute at scale. For our out-of-distribution evaluation, we use the default CLRS-30 benchmark testing size $n=64$. The models are trained with default CLRS-30 benchmark model parameters (Appendix D) [2].
@@ 3.1 Results for Bellman-Ford
On small graphs, ($n=5$), sampling methods extract a single correct minimum-cost path with perfect accuracy from the NN-predicted distributions (Table 1, Table 2). To validate sampling methods, we also extract solutions from the empirical distributions. Similar results for empirical and predicted distributions suggest the NN predicts a good distribution for $n=5$ graphs. The lack of variety in solutions, with sampling methods always extracting the same solution, is expected for small graphs as there is likely only one minimum-cost path. Diversity increases in the case of $n=64$ graphs, where for each method, sampling a predecessor array five times gives a new solution in nearly all cases (Table 1). For larger graphs, Table 2 shows that sampling methods successfully extract correct minimum-cost paths from the empirical distributions (denoted as $\mathcal{P}$). However, extraction is often inaccurate from the distributions of solutions predicted by our models (denoted $\tilde{\mathcal{P}}$). These results permit us to conclude that our models perform well on small graphs, but have difficulty scaling outside of their training distribution. While comparing our approach with the deterministic algorithmic frameworks of Veličković et al. [2], Ibarz et al. [3] is disadvantageous to our method due to its broader scope, it is worth noting that our BF results are markedly worse than the 99% micro-F1 score in [2, 3]. However, our results are encouraging for future work, as we have demonstrated the feasibility of training neural networks to emulate a version of Bellman-Ford that is more powerful than its deterministic variant.
Table 1: Mean proportions of unique and valid solutions when sampling 5 solutions from empirical ($\mathcal{P}$) and model output ($\tilde{\mathcal{P}}$) distributions for each graph in 5 training runs. Rounded figures.
Bellman-Ford
@@
Greedy ($n=5$) $0.20\pm 0.00$ $1.00\pm 0.00$ $0.20\pm 0.00$ $1.00\pm 0.00$
Beam ($n=5$) $0.20\pm 0.00$ $1.00\pm 0.00$ $0.20\pm 0.00$ $1.00\pm 0.00$
Greedy ($n=64$) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.01$ $0.40\pm 0.19$
Beam ($n=64$) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.54\pm 0.22$
Depth-First Search
@@
Upwards ($n=5$) $0.74\pm 0.04$ $0.32\pm 0.06$ $0.52\pm 0.05$ $0.16\pm 0.07$
AltUpwards ($n=5$) $0.51\pm 0.03$ $0.86\pm 0.02$ $0.73\pm 0.05$ $0.56\pm 0.14$
Upwards ($n=64$) $1.00\pm 0.00$ $0.01\pm 0.02$ $1.00\pm 0.00$ $0.00\pm 0.01$
AltUpwards ($n=64$) $1.00\pm 0.00$ $0.05\pm 0.03$ $1.00\pm 0.00$ $0.01\pm 0.02$
@@ 3.2 Results for Depth-First Search
In most cases, our sampling methods perform worse than simply using the most likely predecessor for all $v\in V(G)$, but have the advantage of recovering multiple solutions (Table 1). On smaller graphs, AltUpwards performs well on the model distributions, returning a valid DFS-tree around half of the time. We only consider necessary conditions, causing up to 40% of trees identified as "correct" to be incorrect for $n=5$ (evaluating this metric for $n=64$ is computationally prohibitive in absence of sufficient correctness conditions). DFS is known to be difficult, with NARs giving correct (single) solutions only around a quarter of the time without pointer reversal [2, 3, 4]. This contextualises our results and demonstrates the promise of our approach compared to both classical algorithms and previous work in neural algorithmic reasoning. In terms of the ability of recovering multiple solutions, Table 1 shows that for both considered graph sizes, many unique solutions are obtained, with better accuracy scores for $\mathcal{P}$ than for $\tilde{\mathcal{P}}$, particularly for large graphs, where accuracy is low in general. In fact, for the out-of-distribution testing, our NAR method for DFS performs significantly worse than the current state of the art [3]. Because our method is non-deterministic, this should not be taken as a conclusively negative result: we would expect our method for returning multiple solutions to perform worse than the results for models giving single solutions in [3], as predicting multiple solutions is a more difficult task with more noise. Furthermore, AltUpwards is generally able to recover more unique and more valid solutions than Upwards, but at a high computational cost. Table 1 shows that for large graphs, both sampling methods yield many unique solutions, but these solutions are hardly ever valid for large graphs.
Table 2: Mean accuracies (0 to 1) for all sampling methods, with standard deviations over 5 runs.
@@ Conclusion
We outline a two-part approach to NAR with multiple solutions: we train models to predict distributions of solutions and stochastically extract solutions from those distributions. We demonstrate feasibility, given a good distribution, but also that NNs do not easily learn to predict good distributions. We attribute this difficulty to a number of factors. First, NAR is difficult: previous work has shown that NNs only achieve up to 40% micro-F1 on deterministic DFS without hint pointer reversal, which improves it to nearly 100% [4]. Second, training a model to fit a distribution of solutions does not directly optimize for correct solutions. Solutions must still be extracted from the distribution via stochastic methods, to which the loss metric does not extend (Figure 7). Improving methods for extracting solutions remains an important algorithmic task. Still, with the promise of sampling multiple solutions from a distribution demonstrated, we believe our modifications of the CLRS benchmark act as proof-of-concept for further work. Our method is applicable to other CLRS-30 algorithms such as Minimum Spanning Tree, Bipartite Matching, and Task Scheduling. Having multiple solutions adds essential flexibility to applications of NAR such as cyberphysical systems and might aid in explaining NAR models.
@@ Appendix A Methodology
Figure 2: Parent Tree encoding of a single DFS solution. The start node, 0, is its own parent, represented by the value 0 at index 0. The value 2 at index 1 indicates that vertex 2 is the parent of vertex 1. Similarly, the value 0 at index 2 indicates that vertex 0 is the parent of vertex 2.
@@ C.2 Validating model output distributions for DFS
Appendix D Training Details
D.1 Model architecture and parameters
To give a good baseline, we use the default CLRS Benchmark NAR neural network [2]. Our processor network is an MPNN [10] with a two standard NAR adjustments: namely, gating and triplet reasoning [3]. Our encoder is standard: a linear layer initialised with xavier on scalars. Our decoder is 4 linear layers — three with hidden dimension (128) and a final layer with dimension one. The network is recurrent.
Parameter Value
Processor triplet_gmpnn
Encoder xavier_on_scalars
Decoder 4 linear layers
#Parameters 391225
#Hidden Units 128
Learning Rate 0.001
Batch Size 1
Training Graph Sizes 4, 7, 11, 13, 16
Training Set Size 1000
Testing Graph Sizes 5, 64
Testing Set Size 10
Table 3: Training parameters
D.2 Time characteristics
We run our experiments using a publicly available compute provider that uses NVIDIA P100 GPUs. Including evaluation of our model every 50 steps, training the NARs for 10,000 steps is very time-inexpensive. We summarize the training times in Table 4.
$n=5$ $n=64$
BF 240sec. 248sec.
DFS 531sec. 550sec.
Table 4: Training times for the conducted experiments in seconds